A consumers cares only about the total amount of beer he drinks. Therefore he considers two 12-ounce cans of beer to be as good as one 24-ounce bottle. Suppose these are the only two goods available to this consumer. The price of a 12-ounce can is $1.00, while the price of a 24-ounce bottle is $3.00 due to higher packaging costs.

a) Write down his budget constraint and a utility function that captures his preferences. Draw his budget constraint and three of his indifference curves.

b) What is his optimal consumption bundle? Explain your reasoning.

c) Fixing the price of a 24-ounce beer at $3.00, what must we make the price of a 12-ounce beer to have the consumer purchase both goods?

d) At the price you determine in c) what is the best choice? Explain.


1
Expert's answer
2021-10-01T12:24:21-0400

a)

12 ounce of beer=X

24 ounce of beer=y

Utility will be

"4=2x+y"

Budget constraint

Price of X=1

Price of y=3

"PX(X)+py(y)=I\\X+3y=I"


b)

Optimal consumption bundle

"MRS=\frac{\frac{\delta u}{\delta x}}{\frac{\delta u}{\delta y}}=2\\\frac{px}{py}=\frac{1}{3}"


C)

"MRS=\frac{px}{py}=2\\py=3\\px=2×3=6"


D)

Best choice will be 8 and 6